# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws (Extended Version)

### v2

@@ III-D Switching Law
Proof.
Consider the connection between two constrained arcs. Assume that $\mathcal{P}_{1}\cap\mathcal{P}_{2}\not=\varnothing$. Arbitrarily take a $p^{*}\in\mathcal{P}_{1}\cap\mathcal{P}_{2}$ into consideration. According to Proposition 2, $\forall t\in\left[t_{0},t_{2}\right]$, $\dot{\boldsymbol{x}}=\left({\boldsymbol{A}}-{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p^{*}}^{\top}\right){\boldsymbol{x}}$ holds. Evidently, $\forall p\in\mathcal{P}_{1}\cup\mathcal{P}_{2}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d_{p}\equiv 0$, and $\forall r\in\left[r_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r}{\boldsymbol{x}}\left(t_{1}\right)\equiv 0$. By Proposition 2, $\forall t\in\left[t_{0},t_{2}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ holds. Therefore, $\mathcal{P}_{1}=\mathcal{P}_{2}$, which contradicts ${\mathcal{S}}_{1}\not={\mathcal{S}}_{2}$. Hence, $\mathcal{P}_{1}\cap\mathcal{P}_{2}=\varnothing$.
@@
Proposition 4 provides a necessary and sufficient condition for the feasibility of adjacent arcs near the connection. Note that the equality constraints at the connection time, i.e., (11), (20), (22), and (24), are allowed linearly dependent. When examining feasibility near the connection time, it is sufficient to consider inequality constraints and all linearly independent equality constraints.
IV Keypoints for the Trajectory’s Feasibility
@@ IV-A Additional Constraints at the End of An Arc
      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<{u_{\mathrm{m}}}$.
@@
      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$. Furthermore, $\exists\hat{r}_{p}\in\left[n\right]$, s.t. $\forall r\in\left[\hat{r}_{p}-1\right]$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$, and $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<0$.
@@
      $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$. Furthermore, $\forall r\in\left[n\right]$, $-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$.
@@
Proof.
Proposition 5.1 holds for the same reason as that of Proposition 4.
Consider Proposition 5.2. Proposition 2 implies that $u=-\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}$. $\exists\varepsilon>0$, s.t. $u\left(t\right)\leq{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$ if and only if (1) $\exists\varepsilon>0$, s.t. $u<{u_{\mathrm{m}}}$ for $\left(t_{0},t_{0}+\varepsilon\right)$, or (2) $\exists\varepsilon>0$, s.t. $u\left(t\right)\equiv{u_{\mathrm{m}}}$ on $\left[t_{0},t_{0}+\varepsilon\right]$. By Proposition 2, (1) is equivalent to $\exists\hat{r}_{p}\in\left[0,n\right]\cap{\mathbb{N}}$, s.t. $\forall r\in\left[0,\hat{r}_{p}-1\right]\cap{\mathbb{N}}$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$, and $\frac{\mathrm{d}^{\hat{r}_{p}}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{\hat{r}_{p}}}\left(t_{0}\right)<0$. (2) is equivalent to $\forall r\in\left[0,n\right]\cap{\mathbb{N}}$, $\frac{\mathrm{d}^{r}\left(u-{u_{\mathrm{m}}}\right)}{\mathrm{d}t^{r}}\left(t_{0}\right)=0$. Hence, Proposition 5.2 holds. ∎
@@
Consider two arcs ${\mathcal{S}}_{i}$ with active constraints $\mathcal{P}_{i}$ for $t\in\left[t_{i-1},t_{i}\right]$, $i\in\left\{1,2\right\}$. $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ induce the equality constraints ${\boldsymbol{F}}{\boldsymbol{x}}\left(t_{1}\right)+{\boldsymbol{g}}={\boldsymbol{0}}$ where ${\boldsymbol{F}}$ has full row rank. By Proposition 5, state constraints except $\mathcal{P}_{1}\cup\mathcal{P}_{2}$ and control constraints induce $\widehat{{\boldsymbol{C}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{d}}}<{\boldsymbol{0}}$ and $\widehat{{\boldsymbol{F}}}{\boldsymbol{x}}\left(t_{1}\right)+\widehat{{\boldsymbol{g}}}={\boldsymbol{0}}$, where $\left[\begin{array}[]{c}{\boldsymbol{F}}\\ \widehat{{\boldsymbol{F}}}\end{array}\right]$ has full row rank. Then, the additional end-constraint of ${\boldsymbol{x}}\left(t_{1}\right)$ is defined as ${\mathcal{E}}=\left(\widehat{{\boldsymbol{C}}},\widehat{{\boldsymbol{d}}},\widehat{{\boldsymbol{F}}},\widehat{{\boldsymbol{g}}}\right)$. Furthermore, if $\widehat{{\boldsymbol{C}}}$ and $\widehat{{\boldsymbol{F}}}$ are empty matrices, then ${\mathcal{E}}$ is called empty.
@@ IV-B Tangent Markers in An Arc
Proof.
${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t_{1}\right)+d=0$. $\exists\varepsilon>0$, s.t. $B_{\varepsilon}\left(t_{1}\right)\subset\left(t_{0},t_{2}\right)$ and $\forall t\in{\left.B_{\varepsilon}\left(t_{1}\right)\middle\backslash\left\{t_{1}\right\}\right.}$, ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}\left(t\right)+d<0$. Note that $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}}{\mathrm{d}t^{r}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)={\boldsymbol{c}}^{\top}\widehat{{\boldsymbol{A}}}^{r-1}\left(\widehat{{\boldsymbol{A}}}{\boldsymbol{x}}+\widehat{{\boldsymbol{b}}}\right)$.
If $\hat{r}$ is odd, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d$ strictly crosses 00 at $t_{1}$, which contradicts the tangent condition. Hence, $\hat{r}$ is even. If $\left.\frac{\mathrm{d}^{\hat{r}}}{\mathrm{d}t^{\hat{r}}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}>0$, then ${\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d$ achieves a strict local minimum at $t_{1}$, which contradicts the maximum condition. Hence, $\left.\frac{\mathrm{d}^{\hat{r}}}{\mathrm{d}t^{\hat{r}}}\left({\boldsymbol{c}}^{\top}{\boldsymbol{x}}+d\right)\right|_{t_{1}}<0$. Therefore, (29) holds.
No matter an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ is constrained or not, ${\mathcal{S}}$ can be tangent to the boundary of the state constraint ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ where $p\not\in\mathcal{P}$. It is noteworthy that ${\mathcal{S}}$ can also be tangent to the boundary of the control constraints $u\leq{u_{\mathrm{m}}}$ and $-u\leq{u_{\mathrm{m}}}$. For this case, ${\mathcal{S}}$ should be constrained. Arbitrarily consider $p\in\mathcal{P}$. According to Proposition 2, ±$\pm u\leq{u_{\mathrm{m}}}$ is equivalent to $\mp\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\leq{u_{\mathrm{m}}}$. In other words, the constraints on $u$ are equivalent to the constraints on ${\boldsymbol{x}}$. Hence, Proposition 6 can apply for both state constraints and control constraints. For convenience, $\forall p\in\left[P\right]$, the $p$-th constraint refers to ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$. The $\left(P+1\right)$-th and $\left(P+2\right)$-th one refer to $u\leq{u_{\mathrm{m}}}$ and $-u\leq{u_{\mathrm{m}}}$, respectively.
@@ IV-C Augmented Switching Law
Section III, Section IV-A, and Section IV-B analyze the feasibility of the optimal trajectory at the end points of arcs and the tangent points to the constrained boundaries. To characterize the optimal trajectory, this section proposes the augmented switching law, which represents the input control and the feasibility of the trajectory in a compact form.
@@
Consider a feasible BBS trajectory of problem (1a) with $N$ arcs, denoted as ${\mathcal{S}}_{i}$, $i\in\left[N\right]$. During ${\mathcal{S}}_{i}$, $M_{i}$ tangent markers occur, denoted as ${\mathcal{T}}_{k}^{(i)}$, $k\in\left[M_{i}\right]$. At the connection of ${\mathcal{S}}_{i}$ and ${\mathcal{S}}_{i+1}$, the additional end-constraint is denoted as ${\mathcal{E}}_{i}$. Then, time points at tangent markers and ends of arcs are called keypoints for the trajectory’s feasibility. The augmented switching law (ASL) ${\mathcal{L}}$ is defined in (28).
@@
For convenience of understanding, the chain-of-integrator system with full box state constraints in Example 2 is considered as an example that was discussed in our previous works [20, 32]. The system behaviors where $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$, i.e., unconstrained arcs, are denoted as $\overline{0}$ and $\underline{0}$, respectively. The system behaviors where $x_{k}\equiv x_{{\mathrm{m}}k}$ and $x_{k}\equiv-x_{{\mathrm{m}}k}$, i.e., constrained arcs, are denoted as $\overline{k}$ and $\underline{k}$, respectively. Evidently, during $\overline{k}$ or $\underline{k}$, $u\equiv 0$ and $\forall j\in\left[k-1\right]$, $x_{j}\equiv 0$. The tangent marker where ${\boldsymbol{x}}$ is tangent to $x_{{\mathrm{m}}k}$ and $-x_{{\mathrm{m}}k}$ of even order $2l<k$ in Proposition 6 is denoted as $\left(\overline{k},2l\right)$ and $\left(\underline{k},2l\right)$, respectively. The additional end-constraint =${\mathcal{E}}=\left\{\left(s_{k},r_{k}\right)\right\}_{k=1}^{K}$ means that the additional state constraints $s_{k}$ is active with order $r_{k}$ in Proposition 5, where $s_{k}=\overline{j}$ and $s_{k}=\underline{j}$ refer to the constraints $x_{j}\leq x_{{\mathrm{m}}j}$ and $-x_{j}\leq x_{{\mathrm{m}}j}$, respectively. Note that the constraints on control are inactive at the ends of constrained arcs since $u\equiv 0$ in a constrained arc. Based on the above simplified notations, an example is provided to illustrate the ASL.
@@ V-A Uniqueness of Time-Optimal Control
Proof.
Therefore, both ${\mathcal{S}}_{1}^{(1)}$ and ${\mathcal{S}}_{1}^{(2)}$ are constrained arcs. Assume that $\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}\not=\varnothing$. Arbitrarily consider $p\in\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}$. $\forall t\in\left(t_{0},\min\left\{t_{1}^{(1)},t_{1}^{(2)}\right\}\right)$, $i=1,2$, $\dot{\boldsymbol{x}}_{i}^{*}\left(t\right)=\left({\boldsymbol{A}}-{\boldsymbol{b}}\widehat{{\boldsymbol{a}}}_{p}^{\top}\right){\boldsymbol{x}}_{i}^{*}\left(t\right)$. ${\boldsymbol{x}}_{1}^{*}\left(t_{0}\right)={\boldsymbol{x}}_{2}^{*}\left(t_{0}\right)$ implies that ${\boldsymbol{x}}_{1}^{*}\left(t\right)\equiv{\boldsymbol{x}}_{2}^{*}\left(t\right)$, which contradicts the definition of $t_{0}$. Therefore, $\mathcal{P}_{1}^{(1)}\cap\mathcal{P}_{1}^{(2)}=\varnothing$. Then, $\exists\varepsilon\in\left(0,\min\left\{t_{1}^{(1)},t_{1}^{(2)}\right\}-t_{0}\right)$, s.t. $\forall t\in\left(t_{0},t_{0}+\varepsilon\right)$, $i=1,2$, $p\not\in\mathcal{P}_{1}^{(i)}$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}_{i}^{*}\left(t\right)+d_{p}<0$. Hence, $\forall p\in\left[P\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}_{3}^{*}\left(t\right)+d_{p}={\boldsymbol{c}}_{p}^{\top}\frac{{\boldsymbol{x}}_{1}^{*}\left(t\right)+u{\boldsymbol{x}}_{2}^{*}\left(t\right)}{2}+d_{p}<0$. By Proposition 2, $u_{3}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)\in\left\{\pm{u_{\mathrm{m}}}\right\}$, which contradicts the definition of $t_{0}$. Therefore, $\forall t\in\left[t_{0},t_{\mathrm{f}}^{*}\right]$, ${\boldsymbol{x}}_{1}^{*}\left(t\right)={\boldsymbol{x}}_{2}^{*}\left(t\right)$; hence, $u_{1}^{*}\left(t\right)=u_{2}^{*}\left(t\right)$ a.e. ∎
@@
According to Theorem 1, if the time-optimal trajectory is perturbed by a sufficiently and finitely small perturbation, then the resulting perturbed trajectory achieves a strictly shorter terminal time than the optimal terminal time. The above observation will be applied in Section V-B.

### v3

@@ III-C Switching Law
Proposition 4 provides a necessary and sufficient condition for the feasibility of adjacent arcs near the connection. Note that the equality constraints at the connection time, i.e., (10), (17), (19), and (21), are allowed to be linearly dependent. When examining feasibility near the connection time, it is sufficient to consider inequality constraints and all linearly independent equality constraints.
IV Keypoints for Trajectory Feasibility
@@ IV-A Additional Constraints at the End of An Arc
      $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<{u_{\mathrm{m}}}$.
@@
      $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$. Furthermore, $\exists\hat{r}_{p}\in\left[n\right]$, s.t. $\forall r\in\left[\hat{r}_{p}-1\right]$, $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$, and $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{\hat{r}_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)<0$.
@@
      $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)={u_{\mathrm{m}}}$. Furthermore, $\forall r\in\left[n\right]$, $\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}\widehat{{\boldsymbol{A}}}^{r}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}{\boldsymbol{x}}\left(t_{0}\right)=0$.
@@
[…]
@@ IV-B Tangent Markers in An Arc
Regardless of whether an arc ${\mathcal{S}}=\left(\widehat{{\boldsymbol{A}}},\widehat{{\boldsymbol{b}}},{\boldsymbol{F}},{\boldsymbol{g}},\mathcal{P}\right)$ is constrained, $\mathcal{S}$} can be tangent to the boundary of the state constraint ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ where $p\not\in\mathcal{P}$. It is noteworthy that $\mathcal{S}$} can also be tangent to the boundary of the control constraints $u\leq{u_{\mathrm{m}}}$ and $u\leq{u_{\mathrm{m}}}$. For this case, $\mathcal{S}$} should be constrained. Arbitrarily consider $p\in\mathcal{P}$. According to Proposition 2, $\pm u\leq{u_{\mathrm{m}}}$ is equivalent to $\mp\frac{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}}{\boldsymbol{x}}}{{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}^{r_{p}-1}{\boldsymbol{b}}}\leq{u_{\mathrm{m}}}$. In other words, the constraints on $u$ are equivalent to the constraints on $\boldsymbol{x}$}. Hence, Proposition 6 can apply for both state constraints and control constraints. For convenience, $\forall p\in\left[P\right]$, the $p$-th constraint refers to ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$. The $\left(P+1\right)$-th and $\left(P+2\right)$-th one refer to $u\leq{u_{\mathrm{m}}}$ and $u\leq{u_{\mathrm{m}}}$, respectively.
@@ IV-C Augmented Switching Law
Sections III, IV-A, and IV-B analyze the feasibility of the optimal trajectory at the end points of arcs and the tangent points to the constrained boundaries. To characterize the optimal trajectory, this section proposes the augmented switching law, which represents the input control and the feasibility of the trajectory in a compact form.
@@
Consider a feasible BBS trajectory of problem (1a) with $N$ arcs, denoted as ${\mathcal{S}}_{i}$, $i\in\left[N\right]$. During ${\mathcal{S}}_{i}$, $M_{i}$ tangent markers occur, denoted as ${\mathcal{T}}_{k}^{(i)}$, $k\in\left[M_{i}\right]$. At the connection of ${\mathcal{S}}_{i}$ and ${\mathcal{S}}_{i+1}$, the additional end-constraint is denoted as ${\mathcal{E}}_{i}$. Then, time points at tangent markers and ends of arcs are called keypoints for trajectory feasibility. The augmented switching law (ASL) $\mathcal{L}$} is defined in (22).
@@
For convenience of understanding, the chain-of-integrator system with full box state constraints in Example 2 is considered as an example that was discussed in our previous works [20, 31]. The system behaviors where $u\equiv{u_{\mathrm{m}}}$ and $u\equiv-{u_{\mathrm{m}}}$, i.e., unconstrained arcs, are denoted as $\overline{0}$ and $\underline{0}$, respectively. The system behaviors where $x_{k}\equiv x_{{\mathrm{m}}k}$ and $x_{k}\equiv-x_{{\mathrm{m}}k}$, i.e., constrained arcs, are denoted as $\overline{k}$ and $\underline{k}$, respectively. Evidently, during $\overline{k}$ or $\underline{k}$, $u\equiv 0$ and $\forall j\in\left[k-1\right]$, $x_{j}\equiv 0$. The tangent marker where $\boldsymbol{x}$} is tangent to $x_{{\mathrm{m}}k}$ and $x_{{\mathrm{m}}k}$ of even order $2l<k$ in Proposition 6 is denoted as $\left(\overline{k},2l\right)$ and $\left(\underline{k},2l\right)$, respectively. The additional end-constraint ${\mathcal{E}}=\left\{\left(s_{k},r_{k}\right)\right\}_{k=1}^{K}$ means that the additional state constraints $s_{k}$ is active with order $r_{k}$ in Proposition 5, where $s_{k}=\overline{j}$ and $s_{k}=\underline{j}$ refer to the constraints $x_{j}\leq x_{{\mathrm{m}}j}$ and $x_{j}\leq x_{{\mathrm{m}}j}$, respectively. Note that the constraints on control are inactive at the ends of constrained arcs since $u\equiv 0$ in a constrained arc. Based on the above simplified notations, an example is provided to illustrate the ASL.
@@ V-A Uniqueness of Time-Optimal Control
According to Theorem 1, if the time-optimal trajectory is perturbed by a sufficiently and finitely small perturbation, then the resulting perturbed trajectory achieves a strictly longer terminal time than the optimal terminal time. The above observation will be applied in Section V-B.